OpenStudy (anonymous):

f(x)=(2x)/(x-1) whats the domain

OpenStudy (anonymous):

f(x)= 2x^2/(x^4+1) In regards to 27c, d, e, and f. c>> if f(x)=-1, what is x? My understanding... -1=2x^2/x^4+1 Cross-multiply to get: -x^4-1 = 2x^2 Rearrange: x^4 + 2x^2 + 1 = 0 Factor: (x^2+1)(x^2+1) = 0 x = +i or x = -i

OpenStudy (anonymous):

was that helpful?

OpenStudy (anonymous):

domain is all real values except x=1